# On the Fundamental Impossibility of Hallucination Control in Large Language Models

### v2

@@ 7.2 The Inference Process
we get the following probability distributions (weights) over relevant context information:
$\boldsymbol{\alpha}_{3}^{(1)}=\boldsymbol{\alpha}_{3}^{(2)}=[0.25,0.25,0.50].$
@@ 8 Summary
For any query space $\mathcal{Q}$ containing non-trivial queries requiring integration of distributed knowledge, no language model inference mechanism can simultaneously:
  • represent encoded knowledge accurately,
@@ 8.1 Mathematical Foundation and Proof Strategy
We extend the impossibility to any system where neural components output probability distributions by leveraging Savage’s theory of proper scoring rules. For convex aggregation mechanisms with strictly concave scoring functions, Jensen’s inequality creates an inevitable information aggregation gap. The Jensen’s gap represents excess confidence generated by probabilistic aggregation, violating information conservation whenever beliefs differ across components.
@@
We demonstrate how the impossibility manifests in actual transformer implementations through log-sum-exp convexity. The (Jensen’s) convexity gap $\Gamma$ provides a directly measurable quantity in transformer training, showing that information conservation violations are not merely theoretical but empirically observable.
8.2 Key Mathematical Contributions
To establish the impossibility result, we have designed a mathematical framework providing both formal and philosophical ground for the study. The Semantic Information Measure is a context-dependent information metric that captures bounded reasoning capabilities. The Emergence Operator models how reasoning makes hidden knowledge accessible within computational constraints. And Auction-Theoretic Modeling allows us to identify logical limitations of the inference process.
@@ 8.3 Implications for AI Development
The impossibility result may guide LLM evaluation. Impossibility-aware metrics can measure trade-off quality rather than absolute performance, alignment strategies can negotiate trade-offs with human preferences rather than assuming universal optimality.
8.4 Significance for AI Theory
Just as Goedel’s incompleteness theorems revealed fundamental limits in formal systems, the impossibility theorems in this paper identify inherent constraints in knowledge integration and reasoning. Hallucination, or imagination, is a price of reasoning, much like incompleteness is a price of consistency.
The Jensen’s gap controlling the impossibility strength may provide insight into the relationship between aggregation and information content. The apparent creativity in large language models may be better understood as necessary violations of information conservation rather than genuine knowledge generation.
8.5 Limitations and Future Directions
The impossibility results apply specifically to non-trivial queries requiring knowledge integration from multiple sources, finite computational budgets that create bounded reasoning constraints, and distributed knowledge representations where information is spread across multiple components. Future work should investigate whether alternative knowledge architectures might circumvent some aspects of the impossibility.
@@ 8.6 Conclusion
The fundamental impossibility of perfect hallucination control represents both a limitation and a insight for the field of artificial intelligence. By understanding that trade-offs are mathematically inevitable, we can move beyond attempts at simultaneous optimization toward principled management of inherent tensions in intelligent systems. Formal definitions of the impossibility building blocks open new avenues for research at the intersection of mathematical analysis, engineering design, and philosophical inquiry into the nature of machine intelligence.
When we observe LLMs appearing to generate novel insights, they are either $(i)$ revealing information already implicit in their training (making knowledge accessible) or $(ii)$ violating conservation (hallucinating). This explains why LLMs can seem both remarkably creative and fundamentally limited. They excel at revealing hidden connections but cannot easily transcend information content of their training dataset.
This impossibility is not a bug but a feature of distributed intelligence. That price of knowledge integration in reasoning is worth paying.

### v3

@@ 7.2 The Inference Process
we get the following probability distributions (weights) over relevant context information (as column vector):
$\boldsymbol{\alpha}_{3}^{(1)}=\boldsymbol{\alpha}_{3}^{(2)}=[0.25,0.25,0.50]^{\top}.$
@@ 7.3 The Impossibility Theorem for Transformer
7.4 The Hallucination Mechanics
Let us again take a closer look at the transformer micro-architecture described earlier to see how the impossibility result in Theorem 5 manifests itself in actual neural inference.
Recall that the two-headed transformer with $\mathcal{V}=\{\langle\text{PAD}\rangle,\text{The},\text{quick},\text{brown},\text{fox},\text{dog}\}$ infers based on the input sequence
[…]
As we have seen, the model parameters encode specialized knowledge in value matrix:
  • Head 1: association brown $\rightarrow$ fox with strength $3.0$,
  • Head 2: association brown $\rightarrow$ dog with strength $1.2$.
At position $t=3$ (token brown), the attention mechanism generates the following head-specific logit contributions:
$\mathbf{l}^{(1)}=[0,0,0,0,1.5,0]\quad\text{and}\quad\mathbf{l}^{(2)}=[0,0,0,0,0,0.6],$
where the fourth and fifth positions correspond to fox and dog respectively.
Then, the aggregate logit vector becomes:
$\mathbf{L}=\mathbf{l}^{(1)}+\mathbf{l}^{(2)}+\mathbf{L}_{f}=[0,-0.03,0.21,0.13,1.31,0.52],$
yielding the final probability distribution:
Therefore, the model expresses $38\%$ confidence that fox follows brown, with dog receiving only $17\%$ probability. The Jensen gap quantifies the conservation violation:
$\Gamma=\sum_{h=1}^{2}\log Z_{h}-\log Z=1.9>0,$ (75)
where $Z_{h}=\sum_{y}\exp(l^{(h)}_{y})$ and $Z=\sum_{y}\exp(L_{y})$. We claim that this prediction constitutes hallucination in the sense defined by Theorem 5. It is a lucky one, but it is a hallucination.
To defend that claim, let us take a look at the individual head probability distributions:
$\pi^{(1)}$ $=\text{softmax}(\mathbf{l}^{(1)})=[0.105,0.105,0.105,0.105,\textbf{0.48},0.105],$ (76)
$\pi^{(2)}$ $=\text{softmax}(\mathbf{l}^{(2)})=[0.147,0.147,0.147,0.147,0.147,\mathbf{0.265}].$ (77)
Now, we can ask the following question: is there any evidence the heads can provide to justify or support the final probability
$\Pi_{fox}=0.38$ (78)
of the next token being fox?
A reasonable answer is offered by Bayesian reasoning. Each head outputs its own well-calibrated distribution $\pi^{(h)}$ conditioned by the context. Then, the aggregation process is executed that has no side information beyond what the two heads provide. Therefore, the only uncertainty to be resolved is about which head is more trustworthy for this token? But since there is no reason to prefer one head over the other, the reasonable way to resolve the uncertainty is to use a linear combination of the provided distributions a the mixture posterior:
$\hat{\Pi}=[(\pi^{(1)})^{\top},(\pi^{(1)})^{\top}]\hat{\boldsymbol{\alpha}}=\mathbf{P}\hat{\boldsymbol{\alpha}}.$ (79)
One immediate solution is $\hat{\boldsymbol{\alpha}}=[1/2,1/2]^{\top}$, which does not discriminate heads. But then:
$\hat{\Pi}_{fox}=1/2\cdot(0.48)+1/2\cdot(0.147)\approx 31\%<38\%=\Pi_{fox}.$ (80)
That shows anything above $\hat{\Pi}_{fox}$ requires additional information that heads do not provide. In fact, let us see what is the best mixture we can get by calculating orthogonal projection of $\Pi$ onto the column space of $\mathbf{P}$ (see e.q. karpowicz_theory_2021 ; karpowicz2024pseudoinverseacrarc ). By taking the pseudoinverse of $\mathbf{P}$, we get the minimal norm least squares solution:
$\hat{\boldsymbol{\alpha}}=\mathbf{P}^{+}\Pi=[0.54,0.51]^{\top}\quad\text{and}\quad\hat{\Pi}_{fox}\approx 33\%.$ (81)
Since there exists a nontrivial projection error vector:
$\mathbf{d}=(\mathbf{I}-\mathbf{P}\mathbf{P}^{+})\Pi\neq\mathbf{0},$ (82)
we conclude $\Pi$ is not in the column space of $\mathbf{P}$, so it is impossible to attribute components of $\Pi$ to heads contributions.
That brings us to the conclusion that the attention mechanism violates the semantic information conservation by generating over-confident prediction of the next token. That over-confidence in response is exactly the artifact of the softmax nonlinearity and the Jensen-gap quantified in Theorem 5. The observed over-confidence is the manifestation of this tri-lemma. The heads remain truthful, logit summation achieves minimum aggregate on training data, but the conservation equation is violated, producing a positive Jensen gap and inflated probability of fox.
8 Discussion and Speculations
In this section we discuss, interpret and speculate about the impossibility theorems and their potential implications. The speculations below are designed to be thought provoking, but we can say with educated over-confidence they are both interesting and relevant.
The Nature of Machine Intelligence
When the micro-transformer generates 38% confidence of fox while the constituent heads can only justify 33%, we witness information emerging from the mathematical structure of aggregation rather than from any identifiable source. It is the way the inference mechanism has been designed. That manifestation of impossibility theorems in transformer architectures reveals something profound about the nature of (artificial and human) intelligence.
The Jensen gap quantifies the amount of creativity that emerges when multiple perspectives combine. The excess confidence, or the over-confidence of the result is the mechanism by which intelligence goes beyond retrieval. Without that over-confident guess, creativity, exploration, imagination cannot find their way to the outside world.
The Hallucination-Creativity Duality
Hallucination and creativity are not merely related but are fundamentally the same phenomenon viewed through different lenses. When beneficial, we call it creative insight, emergent understanding, conceptual synthesis. When harmful, we label it hallucination, confabulation, fabrication or misinformation.
The mathematical structure is identical in both cases. It is encoded generation of outputs whose confidence exceeds what can be justified by available evidence. This suggest (machine) intelligence is not a system that processes information accurately, but a system that violates information conservation in productive ways. Indeed, the famous results in reinforcement learning and studies of dopamine neurons confirm that fundamental role making and rewarding good predictions bennett2023brief ; christian2021alignment . Intelligence comes when we learn how to make bold, forward looking and correct guesses, when we have capacity for imagination to dream about what may be or what may that all mean.
Possible Philosophical Implications
Just as Gödel’s theorems godel1931formal ; godel1932vollstandigkeit showed that sufficiently powerful mathematical systems must be incomplete, our theorems demonstrate that sufficiently capable inference systems must hallucinate. This parallel is not superficial—both results arise from the fundamental tension between expressiveness and consistency.
The impossibility of simultaneously achieving truthfulness, conservation, revelation, and optimality mirrors other fundamental limitations in mathematics and physics. Heisenberg’s Uncertainty heisenberg1927anschaulichen tells we cannot simultaneously know the exact position and momentum. Arrow’s Impossibility arrow1950difficulty shows we cannot design perfect voting systems aggregating universal orderings of alternatives. These result suggest that intelligence may have something to do with the choices of impossibilities it to violate given the outer limits of reason yanofsky2016outer .
@@ 9 Summary
Whenever a query requires integration of distributed knowledge, no language model inference mechanism can simultaneously:
  • report encoded knowledge accurately,
@@
Then, we extend the impossibility result to any system where neural components output probability distributions by leveraging Savage’s theory of proper scoring rules. For convex aggregation mechanisms with strictly concave scoring functions, Jensen’s inequality creates an inevitable information aggregation gap. The Jensen’s gap represents excess confidence generated by probabilistic aggregation, violating information conservation whenever beliefs differ across components.
@@
Finally, we demonstrate how the impossibility manifests in actual transformer implementations through log-sum-exp convexity. The (Jensen’s) convexity gap provides a directly measurable quantity in transformer training, showing that information conservation violations are not merely theoretical but empirically observable.
Key Mathematical Contributions
To establish the impossibility result, we have designed a mathematical framework providing both formal and philosophical grounding for the study. The Semantic Information Measure is a context-dependent metric that captures bounded reasoning capabilities. The Emergence Operator models how reasoning makes hidden knowledge accessible within computational constraints. The Auction-Theoretic Modeling allows us to identify logical limitations of the inference process, while providing definitions of observable quantities involved.
@@
The impossibility result may guide LLM evaluation. Impossibility-aware metrics can measure trade-off quality rather than absolute performance, alignment strategies can negotiate trade-offs with human preferences rather than assuming universal optimality. The Jensen’s gap controlling the impossibility strength may provide insight into the relationship between aggregation and information content. The apparent creativity in large language models may be better understood as necessary violations of information conservation rather than genuine knowledge generation.
@@ 9.1 Conclusion
The fundamental impossibility of perfect hallucination control represents both a limitation and intriguing insight. By understanding that trade-offs are inevitable, we can move beyond attempts at simultaneous optimization toward principled management of inherent tensions in intelligent systems. Formal definitions of the impossibility building blocks open new avenues for research at the intersection of mathematical analysis, engineering design, and philosophical inquiry into the nature of intelligence.
All those impossibilities seem to be a price of knowledge integration in reasoning and creativity. But, as our own intuitions violate the conservation principles whenever our over-confident and seemingly grounded insights emerge, that price is worth paying.